OpenStudy (anonymous):

If the instructions for a problem ask you to use the smallest possible domain to completely graph two periods of y = 5 + 3 cos 2(x -pi/3 ), what should be used for Xmin and Xmax? Explain your answer.

OpenStudy (anonymous):

@jim_thompson5910 can you help me on this one too?

jimthompson5910 (jim_thompson5910):

what is the period? what is the phase shift?

jimthompson5910 (jim_thompson5910):

If you're not sure, then in general if y = A*cos(Bx - C) + D then the period is T = 2pi/B and the phase shift is C/B

OpenStudy (anonymous):

@jim_thompson5910 what do you do after figutingout period anf shift??

OpenStudy (anonymous):

@wio how do you do this??

jimthompson5910 (jim_thompson5910):

what did you get for the period and phase shift?

OpenStudy (anonymous):

i got a pi for period

jimthompson5910 (jim_thompson5910):

B = 1 T = 2pi/B T = 2pi/1 T = 2pi

jimthompson5910 (jim_thompson5910):

so the period is 2pi

jimthompson5910 (jim_thompson5910):

that means that your window must be 2pi units wide to see a full period (from start to end)

OpenStudy (anonymous):

ohh yea

jimthompson5910 (jim_thompson5910):

now what's your phase shift?

OpenStudy (anonymous):

3/1 sp 3??

OpenStudy (anonymous):

*so

jimthompson5910 (jim_thompson5910):

w = phase shift w = C/B w = (pi/3)/1 w = pi/3

jimthompson5910 (jim_thompson5910):

this means that the graph of cos(x) has been shifted to the right pi/3 units

OpenStudy (anonymous):

ah-ha!!

jimthompson5910 (jim_thompson5910):

so you can start at pi/3 and go 2pi units to get to pi/3 + 2pi = pi/3 + 6pi/3 = 7pi/3

jimthompson5910 (jim_thompson5910):

making xmin = pi/3 xmax = 7pi/3 that will give you a window that's wide enough to see exactly one period

OpenStudy (anonymous):

ohhhh okay!!! thank you so much!!!

jimthompson5910 (jim_thompson5910):

np